Solve the system of equations by graphing.
y-2x=0
y=4x-1
What is the solution of the system of equations?

Answer by fractalier(2101)

From
y - 2x = 0
y = 4x - 1
I can show you the solution but graphing it here is impossible...
Normally you can set up an x-and-y chart and plug in values to get the points you plot...or you can put each line in slope-intercept form and graph...
In either case, the place where the lines intersect is your solution...
In this case, this is x = 1/2 and y = 1 or (1/2, 1)